OpenStudy (anonymous):

WHAT IS THE ECCENTRICITY OF THE GRAPH OF [(X-2)^2/121]+[(Y+6)^2/64]=1?

OpenStudy (cwrw238):

b^2 = a^2( 1 - e^2) for x^1 / a^2 + y^2 / b^2 = 1

OpenStudy (anonymous):

huh?

OpenStudy (cwrw238):

compare the denominators in the question with the above a^2 = 121 and b^2 = 64 e = eccentricity

OpenStudy (cwrw238):

so substituting in the formula 64 = 121(1 - e^2) 121e^2 = 121 - 64 e^2 = 57 / 121 can you finish this?

OpenStudy (anonymous):

yea i got it thanks

OpenStudy (cwrw238):

yw